# 6G: The Intelligent Network of Everything – A Comprehensive Vision, Survey, and Tutorial

### v2

@@ I-D7 Activities in 2024
In May, the South Korean mobile operator KT Corporation and Nokia announced their partnership in developing O-RAN technologies for 6G networks. The National Telecommunications and Information Administration of the US released a request for comment on 6G, asking for opinions of experts from academia and industry on the timing of trials/commercialization, societal benefits, and disaster resilience of 6G. A Japanese consortium, NTT Docomo, NICT, Panasonic, and SKY Perfect JSAT Corporation, informed their successful field test in the 38 GHz spectrum band to simulate the usage of high-altitude platform stations (HAPSs) at an altitude of 4 kilometers using a small plane Cessna. The test was claimed to be the world’s first in kind, providing practical progress toward the use of NTNs at high altitudes in 6G. In June, Nokia and the Gati Shakti Vishwavidyalaya university announced that they have formed a partnership to conduct 5G/6G research on transportation and smart factory use cases. Nokia and Nordic mobile operator Telia reported their successful field trial at the upper 6 GHz spectrum, agreed on at the WRC 2023. These results pave the way toward the needed capacity enhancements in the 6G era using the mid-band spectrum.
I-E Literature Review on 6G Visions and Surveys
This section reviews the main 6G literature in terms of vision and survey articles.
I-E1 6G Vision Articles
@@ I-E Literature Review on 6G Visions and Surveys
In 2021-2024, a few 6G visions were introduced each year. In [61], the authors’ vision was based on four new paradigm shifts, including integrated space-air-ground networks, spectrum operation at sub-6 GHz, mmWave, THz, and optical frequencies, AI-empowered mobile networks, and comprehensive network security. In [62], the vision of the 6G ecosystem was built on the current research status and anticipated future trends. The article [63] studied broadband connectivity for 6G to score the target data rate of 1 Tbit/s. The authors introduced a comprehensive set of technological enablers at the spectrum, infrastructure, and algorithmic levels. In [64], a top-down vision of the 6G ecosystem was provided, discussing lifestyle and societal changes, applications and their technical requirements, key challenges and opportunities for all layers, new spectrum bands and deployment scenarios, design principles and required changes in radio access and core network architectures, novel PHY layer solutions, propagation characteristics, real-time signal processing, and radio frequency (RF) transceiver design. The work [65] introduced the 6G vision of the EU’s 6G flagship program Hexa-X. The five main contributions include a joint academic and industry perspective of 6G; review of concurrent 6G initiatives; identification of six key challenges; insightful discussion on application scenarios; and analysis of technological transformations in the radio, network, and orchestration domains.
In [66], the proposed 6G vision focused on immersive, intelligent, and ubiquitous applications and the needed technologies. The paper [67] introduced its 6G exploration with a discussion on the evolution of mobile networks, key features and the required capabilities, architectural perspectives, hardware-software evolution, European 6G projects, and main applications and their challenges. A vision of 6G hyper-connectivity, with nearly unlimited data rates, coverage, and computing, was given in [68]. In [69], the authors introduced a value-oriented 6G by defining multi-dimensional performance indicators, enabling technological elements, and a case study on intelligent multiple access. The work [70] briefly presented the 6G vision of the NGMN Alliance, focusing on the key use cases. The paper [71] introduced an evolutionary framework for 6G. The aspects explored are ITU-R activities, usage scenarios, required capabilities, and potential technologies. In [72], the authors focused on building a 6G vision based on four use cases, i.e., Internet of Senses, connected intelligent machines, digitalized and programmable physical world, and connected sustainable world. In [73], the European 6G vision was introduced, based on the perspective of the EU’s collaborative 6G research program SNS JU.
I-E2 6G Survey Articles
@@
The authors of [89] identified 12 scientific challenges for 6G to rethink the theoretical foundations of traditional communication and to tackle the new paradigm of the convergence of communication and beyond-communication technologies. The work [90] examined 6G architectures and technologies to form an enabling platform for innovative future applications. The survey [91] built a comprehensive picture of 6G by analyzing recent advances, key enablers, and technological challenges. In [92], the authors provided an insightful analysis of 6G enabling technologies and discussed the design of a comprehensive 6G network architecture. The article [93] reviewed the key technologies of 6G networks, focusing on the extreme performance that they can provide in diverse dimensions. In [94], 6G evolution was discussed, with a special focus on the usage scenarios of IMT-2030. The work [95] explored the evolution from 5G-Advanced to 6G from the perspective of three novel services, including immersive communications, everything connected, and high-accuracy sensing. In[96], the authors designed and reviewed an end-to-end 6G system, based on the views of the EU’s 6G flagship project Hexa-X-II.
Figure 5: Content of the article.
I-F Contributions of the Article
In this section, the contributions of this article are summarized. For clarity, they are presented in the order of their appearance in the article.
  • Comprehensive Vision, Survey, and Tutorial on 6G: The main contribution of this article is to provide an insightful vision, contemporary survey, and instructive tutorial for 6G, all in the same package. To the best of our knowledge, this is one of the most comprehensive 6G articles written in the literature.
  • Evolution from 1G to 6G: We provide an insightful introduction to the evolution of mobile communications by presenting the essentials of each generation. In particular, we discuss how disruptive different generations have been and envision what will be the next major disruption enabled by 6G.
  • Overview of 5G: We present a compact overview of 5G. The focus is on 3GPP’s 5G System and Releases 15, 16, 17, 18, and 19.
  • Development Process of 6G: We describe the development process of 6G at a high level. Specifically, we discuss the roles of academia, industry, ITU-R, and 3GPP. The estimated timeline is also introduced.
  • Research Activities toward 6G: We review the main worldwide 6G research activities in a chronological order (2018-2024). Various types of activities are considered, such as projects, collaborations, funding, standardization, experiments, events, and publications.
  • Literature Review on 6G Visions and Surveys: We review the main 6G vision and survey articles.
  • 6G Vision: We introduce an insightful and coherent 6G vision. In our vision, we identify the big picture, fundamental elements, disruptive applications, key use cases, target requirements, potential technologies, and defining features. Our vision is well aligned with ITU-R’s ”IMT-2030 Framework”, but more far-reaching.
  • Fundamental Elements of 6G: We identify and discuss three fundamental elements of 6G, i.e., wireless, AI, and IoE.
  • Disruptive Applications for 6G: We define 12 disruptive applications for 6G in three categories, i.e., human-machine interactions, smart environments, and connected autonomous systems.
  • Key Use Cases for 6G: We identify eight primary use cases in two different categories for 6G, i.e., five communication-oriented and three beyond-communication-oriented.
  • Main Performance Requirements for 6G: We review the main performance requirements of IMT-2030, introduced by ITU-R, and compare them to the typically proposed values in the 6G literature.
  • AI/ML for 6G: We provide a brief introduction to the concepts of AI and ML, and discuss three promising ML methods for 6G.
  • Potential Technologies for 6G: We identify a comprehensive set of potential 6G technologies. Each technology is reviewed in a tutorial manner, along with a survey of the main literature. Specifically, we discuss the vision, introduction, past, present, opportunities, challenges, literature, and future directions. To the best of our knowledge, this is the most comprehensive set of 6G technologies reviewed in the literature.
  • Defining Features for 6G: We specify 12 main features that define the essence of 6G.
  • 6G in a Nutshell: We provide a compact summary of our 6G vision in bullet points so that the readers can obtain the big picture of 6G at one glance.
  • 7G Vision: To place our 6G vision in a broader perspective, we discuss what comes after 6G and sketch the first high-level vision on 7G. The focus is on the main disruption, fundamental elements, and possible applications of 7G. To the best of our knowledge, this is the first 7G vision in the literature.
The remainder of this paper is organized as follows. In Section II, our 6G vision is introduced. Three fundamental elements of 6G are identified in Section III. Section IV defines a comprehensive set of disruptive 6G applications. Key use cases are proposed for 6G in Section V. Section VI reviews the main performance requirements that 6G is expected to satisfy. AI/ML is discussed for 6G in Section VII. In Section VIII, a comprehensive set of potential 6G technologies is reviewed. Twelve features that define the essence of 6G are identified in Section IX. Section X summarizes our 6G vision in the bullet points. In Section XI, the post-6G era is speculated and a high-level vision is sketched for 7G. Finally, conclusions are drawn in Section XII. The list of abbreviations can be found in Appendix. The outline of the article is summarized in Figure 5.

### v3

@@ I-D7 ACTIVITIES IN 2024
In May, the South Korean mobile operator KT Corporation and Nokia announced their partnership in developing O-RAN technologies for 6G networks. Ericsson published a white paper on 6G spectrum considerations, discussing the importance of mid and sub-THz bands to enable mobile life in the 2030s [96]. The National Telecommunications and Information Administration of the US released a request for comment on 6G, asking for opinions of experts from academia and industry on the timing of trials/commercialization, societal benefits, and disaster resilience of 6G. A Japanese consortium, NTT Docomo, NICT, Panasonic, and SKY Perfect JSAT Corporation, informed their successful field test in the 38 GHz spectrum band to simulate the usage of high-altitude platform stations (HAPSs) at an altitude of 4 kilometers using a small plane Cessna. The test was claimed to be the first in kind, providing practical progress toward the use of NTNs at high altitudes in 6G. In June, Nokia and the Gati Shakti Vishwavidyalaya university announced that they have formed a partnership to conduct 5G/6G research on transportation and smart factory use cases. Nokia and Nordic mobile operator Telia reported their successful field trial at the upper 6 GHz spectrum, agreed on at the WRC 2023. These results pave the way toward the needed capacity enhancements in the 6G era using the mid-band spectrum.
TABLE II: Comparison of comprehensive 6G surveys
Ref. Evolution 1G-6G 5G Overview Develop. Process Activities Vision Appli-cations Use Cases Require-ments AI/ML Techno-logies 7G Vision
Alsharif -20 [58] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Akyildiz -20 [61] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Khan -20 [62] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Bariah -20 [64] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Akhtar -20 [69] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
You -21 [77] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Rajatheva -21 [79] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Jiang -21 [82] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Bhat -21 [78] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Alwis -21 [83] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Tataria -21 [80] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Alsabah -21 [85] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Uusitalo -21 [76] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Alraih -22 [92] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Hakeem -22 [91] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Salameh -22 [93] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Wang -22 [90] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Asghar -22 [94] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Akbar -22 [95] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Banafaa -23 [97] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Shen -23 [98] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Wang -23 [99] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Quy -23 [100] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Mohsan -23 [101] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Ishteyaq -24 [102] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Shafi -24 [103] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
Giuliano -24 [104] $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
This survey $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$ $\blacksquare$
$\blacksquare$ = EXCEPTIONAL COVERAGE, $\blacksquare$ = HIGH COVERAGE, $\blacksquare$ = MEDIUM COVERAGE, $\blacksquare$ = LOW COVERAGE
I-E CONTRIBUTIONS OF THE ARTICLE
In this section, the contributions of this article are summarized. In Table II, we compare our article against other comprehensive 6G surveys in the literature.
  • Comprehensive Vision, Survey, and Tutorial on 6G: The main contribution of this article is to provide a thorough vision, survey, and tutorial on 6G, all in the same package. This can be seen as a sum of the contributions of all individual sections. To the best of our knowledge, this is the most comprehensive 6G survey in the literature, as indicated in Table II. The contributions of each section are introduced below. They are presented in the order of their potential value to the readers.
    • Potential Technologies for 6G: We identify and review 27 potential 6G technologies in 10 different technology categories. Each technology is reviewed in a tutorial manner, using the same template of vision, introduction, past and present, opportunities and challenges, and literature and future directions. To the best of our knowledge, this is the most comprehensive set of 6G technologies reviewed in a single article, as indicated in Table II. In addition, our tutorial presentation approach is unique in the 6G literature.
    • 6G Vision: We introduce an insightful 6G vision of smart wireless world via 6G-enabled mobile intelligence. In this vision, 6G will evolve toward the Intelligent Network of Everything while serving as a fruitful platform for mobile intelligence to grow toward its potential. Note that our vision is well aligned with the IMT-2030 Framework of ITU-R, although more far-reaching. Based on our 6G vision, we can already see some major trends that evolve beyond 6G and sketch the first high-level vision for 7G. Our insightful and far-reaching 6G vision is novel in the literature.
    • 7G Vision: To place our 6G vision in a broader perspective, we discuss what comes after 6G and sketch a high-level vision for 7G. First, we briefly describe the evolution path from 1G to 7G, with the cycles of two generations. Then, we introduce our 7G vision, focusing on the main disruption, fundamental elements, and possible applications of 7G. Our 7G vision can be summarized in a few words: ”Deeply Intelligent Network of Hyperverses”. To the best of our knowledge, this is the first structured 7G vision in the literature. In addition, the introduced evolution road from 1G to 7G is also a novel view.
    • AI/ML for 6G: Since AI/ML is considered the most revolutionizing technology for 6G, we dedicate a whole section for it. Specifically, we introduce the concepts of AI and ML, review three promising ML methods for 6G, and discuss the opportunities and challenges involved in integrating AI/ML into 6G networks. Most of the existing 6G surveys do not have a dedicated section for 6G AI/ML, with a generic introduction to AI/ML.
    • Disruptive Applications for 6G: We define 12 disruptive applications for 6G in three categories: human-machine interactions, smart environments, and connected autonomous systems. To the best of our knowledge, this is a unique set of categories/applications in the 6G literature.
    • Key Use Cases for 6G: Since 6G is expected to expand its capabilities beyond communication, we identify five communication-oriented and three beyond-communication-oriented use cases. To the best of our knowledge, this is a unique set of categories/use cases in the 6G literature.
    • Main Performance Requirements for 6G: We review the main performance requirements of IMT-2030, introduced by ITU-R, and compare them to the typically proposed values in the 6G literature. To the best of our knowledge, this is the first time when IMT-2030 requirements are compared to the typical values in the 6G literature.
    • Fundamental Elements of 6G: We envision that 6G will be based on three fundamental pillars, i.e., wireless, AI, and IoE. This is a simple and insightful view.
    • Defining Features for 6G: We identify 12 main features that define the essence of 6G: extreme capacity &\&& performance, ultra-flexible &\&& agile, highly intelligent &\&& aware, ubiquitously available &\&& reliable, truly green &\&& sustainable, and thoroughly secure &\&& trustworthy. To the best of our knowledge, this is a unique set of features proposed in the 6G literature.
    • 6G in a Nutshell: We provide a compact summary of our 6G vision in bullet points, allowing readers to grasp the big picture of 6G at one glance. To the best of our knowledge, this style is unique in the 6G literature.
    • Development Process of 6G: We provide an up-to-date and thorough overview of the 6G development process. Specifically, we discuss the roles of ITU-R and 3GPP, as they are the most important standardization bodies in 6G development. The estimated timeline is also introduced. As indicated in Table II, this is the most comprehensive overview of the 6G development process in the literature.
    • Research Activities toward 6G: We review the main worldwide 6G research activities in a chronological order from 2018 to 2024. Various types of activities are considered, such as projects, collaborations, funding, standardization, experiments, events, and publications. This is one of the most thorough overviews of 6G research activities in the literature, as indicated in Table II.
    • Previous Works: We provide a thorough and up-to-date review of the main prior works in terms of 6G visions and surveys, comprising over 60 articles. To the best of our knowledge, this is one of the most comprehensive reviews of the existing 6G visions and surveys in the literature.
    • Overview of 5G: We present an up-to-date overview of 5G in a compact form, focusing on the past, present, and future. In particular, we review the 3GPP 5G System and Releases 15, 16, 17, 18, and 19.
    • Evolution from 1G to 6G: We provide an insightful introduction to the evolution of mobile communications by highlighting the essentials of each generation. In particular, we discuss how disruptive different generations have been and envision what will be the next major disruption enabled by 6G. To the best of our knowledge, this is a novel view on the evolution of mobile networks.
The remainder of this paper is organized as follows. Section II reviews the previous works in the literature. In Section III, our 6G vision is introduced. Three fundamental elements of 6G are identified in Section IV. Section V defines a comprehensive set of disruptive 6G applications. Key use cases are proposed for 6G in Section VI. Section VII reviews the main performance requirements that 6G is expected to satisfy. AI/ML is discussed for 6G in Section VIII. In Section IX, a comprehensive set of potential 6G technologies is reviewed. Twelve features that define the essence of 6G are identified in Section X. Section XI summarizes our 6G vision in the bullet points. In Section XII, the post-6G era is speculated and a high-level vision is sketched for 7G. Finally, conclusions are drawn in Section XIII. The list of abbreviations can be found in Appendix. The outline of the article is summarized in Figure 1.
II PREVIOUS WORKS
In this section, we review the main prior works in terms of 6G vision and survey articles. At the end, we discuss the justification of our work with respect to these previous studies.
II-A 6G VISION ARTICLES
@@ II PREVIOUS WORKS
In 2021-2024, a few 6G visions were introduced each year. In [77], the authors’ vision was based on four new paradigm shifts, including integrated space-air-ground networks, spectrum operation at sub-6 GHz, mmWave, THz, and optical frequencies, AI-empowered mobile networks, and comprehensive network security. In [78], the vision of the 6G ecosystem was built on the current research status and anticipated future trends. The article [79] studied broadband connectivity for 6G to score the target data rate of 1 Tbit/s. The authors introduced a comprehensive set of technological enablers at the spectrum, infrastructure, and algorithmic levels. In [80], a top-down vision of the 6G ecosystem was provided, discussing lifestyle and societal changes, applications and their technical requirements, key challenges and opportunities for all layers, new spectrum bands and deployment scenarios, design principles and required changes in radio access and core network architectures, novel PHY layer solutions, propagation characteristics, real-time signal processing, and radio frequency (RF) transceiver design. The work [76] introduced the 6G vision of the European 6G flagship program Hexa-X. The five main contributions include a joint academic and industry perspective of 6G; review of concurrent 6G initiatives; identification of six key challenges; insightful discussion on application scenarios; and analysis of technological transformations in the radio, network, and orchestration domains.
In [90], the proposed 6G vision focused on immersive, intelligent, and ubiquitous applications and the needed technologies. The paper [91] introduced its 6G exploration with a discussion on the evolution of mobile networks, key features and the required capabilities, architectural perspectives, hardware-software evolution, European 6G projects, and main applications and their challenges. A vision of 6G hyper-connectivity, with nearly unlimited data rates, coverage, and computing, was given in [105]. In [106], the authors introduced a value-oriented 6G by defining multi-dimensional performance indicators, enabling technological elements, and a case study on intelligent multiple access. The work [107] briefly presented the 6G vision of the NGMN Alliance, focusing on the key use cases. The paper [108] introduced an evolutionary framework for 6G. The aspects explored are ITU-R activities, usage scenarios, required capabilities, and potential technologies. In [109], the authors focused on building a 6G vision based on four use cases, i.e., Internet of Senses, connected intelligent machines, digitalized and programmable physical world, and connected sustainable world. In [110], the European 6G vision was introduced, based on the perspective of the collaborative 6G research program SNS JU.
II-B 6G SURVEY ARTICLES
@@
The authors of [113] identified 12 scientific challenges for 6G to rethink the theoretical foundations of traditional communication and to tackle the new paradigm of the convergence of communication and beyond-communication technologies. The work [100] examined 6G architectures and technologies to form an enabling platform for innovative future applications. The survey [101] built a comprehensive picture of 6G by analyzing recent advances, key enablers, and technological challenges. In [102], the authors provided an insightful analysis of 6G enabling technologies and discussed the design of a comprehensive 6G network architecture. The article [103] reviewed the key technologies of 6G networks, focusing on the extreme performance that they can provide in diverse dimensions. In [114], 6G evolution was discussed, with a special focus on the usage scenarios of IMT-2030. The work [104] explored the evolution from 5G-Advanced to 6G from the perspective of three novel services, including immersive communications, everything connected, and high-accuracy sensing. In [115], the authors designed and reviewed an end-to-end 6G system, based on the views of the European 6G flagship project Hexa-X-II.
II-C DISCUSSION
Over the years, 6G has been reviewed from a broad range of aspects in the literature, such as the expected timeline, research activities, novel features, performance requirements, usage scenarios, driving applications, network architectures, and enabling technologies. Throughout the 6G literature, one can notice three common factors above all. First, the performance of 6G is expected to be significantly improved in all dimensions compared to that of 5G. Second, the number and variety of connected objects are expected to grow notable. This is seen as the evolution from IoT to IoE. Third, AI/ML is expected to play a key role in 6G, making it a highly intelligent entity. All these findings from the existing literature support our vision where 6G will be based on extreme wireless, pervasive AI/ML, and ubiquitous IoE, and evolve toward the Intelligent Network of Everything. Due to the years of global 6G research, ITU-R has completed the visioning phase of the IMT-2030 development process and moved to the second phase of defining the performance requirements and evaluation methodology. In the development of technical details, 3GPP is expected to begin the 6G standardization process after mid-2025. At this point, 6G literature is missing an up-to-date overview of the entire 6G research field, providing the latest details in all of the aforementioned 6G aspects. The time is also right to begin visioning beyond 6G. To close this gap in the literature, we provide a contemporary vision, survey, and tutorial on 6G, covering a wide range of timely topics and sketching the first high-level vision beyond 6G.